# How Efficient Are Diffusion Language Models? A Critical Examination of Efficiency Evaluation Practices

### v1

@@ Abstract
Diffusion language models (DLMs) have emerged as a promising alternative to the long-dominant autoregressive (AR) paradigm, offering a parallelable decoding process that could yield greater efficiency. Yet, in practice, current open-source DLMs often underperform their AR counterparts in speed, limiting their real-world utility. This work presents a systematic study of DLM efficiency, identifying key issues in prior evaluation methods. Through empirical benchmarking and a roofline-based theoretical analysis, we demonstrate that AR models generally achieve higher throughput, while DLMs consistently lag. We also investigate acceleration strategies, finding that techniques like dual cache and parallel decoding mainly offer gains at small batch sizes, with their benefits diminishing upon scaling. Our findings underscore the necessity of robust evaluation methods and improved acceleration strategies to advance research on DLMs.
@@ 1 Introduction
Diffusion language models (DLMs) have recently emerged as a competitive alternative to the long-dominant autoregressive (AR) approach [1] for natural language generation. Despite the relatively low profile of non-autoregressive research, its persistent advancement has now yielded performant models like LLaDA [2] that are competitive with their AR counterparts. This represents a significant departure from the sequential generation paradigm that has defined the field for years.
In contrast to AR models, which generate text one token at a time, DLMs produce multiple tokens in parallel via an iterative denoising process. This approach theoretically offers a path to overcome the efficiency bottleneck of sequential generation in AR models. In practice, however, a “paradox” has become increasingly evident: current open-source DLMs often demonstrate slower inference speeds than their AR counterparts of similar scale. For example, LLaMA3-Instruct-8B [3] exhibits an inference throughput 13.7 × greater than that of LLaDA-Instruct-8B on evaluation benchmarks [4]. This “theoretically fast but practically slow” predicament greatly limits the deployment of DLMs in real-world applications, as their potential speed advantages are negated by implementation inefficiencies.
@@
  • Evaluation Scope: Prior evaluations often compare DLMs and AR models in limited conditions, such as single instance inference or fixed generation length. These constrained conditions fail to capture performance across diverse settings and poorly represent real-world use.
@@
  • Throughput modeling and theoretical analysis: We develop a theoretical analysis that decomposes inference throughput into hardware-side compute capacity (FLOPs/s) and model-side compute efficiency (FLOPs/token). Using the roofline model, we analytically characterize the throughput behavior of autoregressive, diffusion, and block diffusion models under varying sequence lengths and batch sizes.
@@ 2.1.3 Block Diffusion Language Models
Block diffusion models combines autoregressive dependencies across blocks with parallel diffusion refinement within each block, maintaining long-range contextual coherence while enabling efficient parallel generation. This architecture also naturally supports KV caching, enhancing inference throughput. It introduces a complementary attention mask that enables bidirectional attention within blocks while preserving autoregressive dependencies across them. During inference, this design allows the model to perform autoregressive generation across blocks and parallel diffusion refinement within each block, leading to a significant reduction in training cost and improved inference efficiency. For example, BD3-LM [5], trained from scratch, demonstrates that block-level diffusion can substantially improve efficiency without sacrificing generation quality. Building on this direction, D2F and Fast-dLLM v2 [6] show that block diffusion-style generation can also be achieved by training from existing models—D2F from diffusion backbones and Fast-dLLM v2 from autoregressive ones.
@@ 2.2.1 Reducing Per-step Computational Overhead
Sequence-level KV Cache. One line of works store and reuse key/value states across decoding steps: The conventional KV cache is designed for strictly autoregressive decoding and does not directly apply to DLMs. To address this, recent work shows that redesigning the decoding schedule allows DLMs to recover much of the efficiency benefit of KV caching. For example, Fast-dLLM [7] introduces a DualCache (prefix + suffix) to improve reuse while bounding quality loss. DPad [8] proposes a suffix-window and distance-decay dropout to restrict attention and thereby limit suffix token caching and computation. And Sparse‑dLLM [9] develops a training-free dynamic cache eviction framework that retains only salient token states in cache and evicts less relevant prefix/suffix entries to improve throughput.
Step-level Feature Reuse. Another line of works reuse stable intermediate representations inside each denoising iteration. dKV-Cache [10] introduces a delayed caching scheme, where a token’s key and value states are cached one denoising step after it is decoded. dLLM-Cache [11] uses feature similarity to identify stable tokens and reuse their cached features., and FreeCache [12] reuses stable KV states of early clean tokens while updating only actively changing ones. These methods are complementary to sequence-level caching.
@@ 2.2.2 Reducing Diffusion Sampling Steps
Step Reduction via Distillation. A key direction for accelerating DLMs is reducing the number of denoising steps through distillation. Early works on diffusion models, such as Progressive Distillation [13], adopt teacher–student schemes that iteratively halve the required sampling steps. Di4C [14] extends this approach to discrete diffusion by explicitly distilling inter-token correlations. Recently, DLM-One [15] apply these distillation ideas to continuous DLMs, using score-based distillation to train a one-step generator that produces the full sequence in a single forward pass, achieving substantial speedups while maintaining near-teacher quality.
@@ 3 Efficiency Evaluation Issues
Efficiency Metrics. Efficiency reporting are inconsistent across studies, making it difficult to compare models fairly. Some studies report latency per token [12], while others focus on throughput (tokens per second) on specific hardware setups [9, 10]. These metrics, while useful, often fail to capture the complete picture. For example, a model might show high throughput, but if each step involves heavy computation, the overall cost could still be high. This is especially relevant for DLMs, which often require more operations per token than AR models. To support clearer and more practical comparisons, it would be helpful to include standardized metrics that consider both computational cost and decoding performance. Examples might include FLOPs per token or throughput under fixed resource constraints. These indicators can offer a more balanced view of model efficiency across different architectures.
@@ 4.1 Overview and Research Questions
RQ2: How can we theoretically explain variations in inference efficiency? We aim to provide a theoretical perspective by decomposing throughput into effective hardware performance (FLOPs per second) and model-specific workload (FLOPs per generated token). This will provide a clear explanation for the empirical results in RQ1, identify key computational bottlenecks, and suggests a pathway for developing future acceleration methods.
@@ 4.2.2 Benchmark and Scenarios
We use the GSM8K dataset for all performance benchmarks. The prompt length is controlled by adjusting the number of examples (i.e., 0-shot and 5-shot) in the prompt. For generation length, DLMs allow direct parameter control. For AR and block diffusion models, we enforce generation to continue until the target length is met (i.e., ranging from 64 to 2048), even after an <eos> token is produced, to ensure a fair comparison of throughput.
@@ 4.2.3 Models and Acceleration Methods
Specifically for the LLaDA model, we evaluate two acceleration strategies introduced in the Fast-dLLM series: the dual cache strategy (to reduce per-step computation) and the confidence-aware parallel decoding strategy (to reduce the number of steps). For the block diffusion, which inherently supports KV cache utilization due to its block-wise autoregressive design, we analyze the confidence-aware parallel decoding strategy. All hyperparameters for these acceleration methods are configured according to their original papers’ settings for GSM8K, and we further verify that the performance remains aligned with the original models.
@@ 4.2.4 Evaluation Metrics
Arithmetic Intensity. To better support the theoretical analysis of DLM efficiency, we adopt arithmetic intensity to quantify the relationship between computation and memory access. It is defined as the ratio of computational volume (FLOPs) to memory access volume (MOPs):
$\text{Arithmetic Intensity}=\frac{\#\text{FLOPs}}{\#\text{MOPs}}.$ (1)
This ratio reflects the average number of computations performed for each byte of memory exchanged (read or write). When a task has low arithmetic intensity, it implies that a large volume of memory access is accompanied by a small amount of computation. Its execution speed will be limited by memory bandwidth, a scenario known as being memory-bound. Conversely, when the arithmetic intensity is high, the computational volume far exceeds the memory access volume. In this case, the performance bottleneck shifts to the GPU’s computational capacity, a scenario known as being compute-bound.
To provide an intuitive understanding of the relationship between arithmetic intensity and performance bottlenecks, we employ the roofline model [16, 17], a framework that connects a system’s attainable compute performance to its memory bandwidth, providing an intuitive way to identify whether a workload is limited by computation or memory access. In this framework, the arithmetic intensity ridge point ($\text{Arithmetic Intensity}_{\text{ridge}}$) defines the boundary between the two regimes (compute-bound and memory-bound) and is calculated as follows:
@@
According to the roofline model, workloads with arithmetic intensity below this ridge point are memory-bound, while those above it are compute-bound. Hence, arithmetic intensity is a key indicator for analyzing and interpreting performance bottlenecks in model inference.
@@ 5.1.1 Effect of Sequence Length
Experimental Settings. We select LLaDA-8B-Instruct, LLaMA3.1-8B-Instruct, and Fast-dLLM v2-7B as the representative models for DLM, AR and block diffusion. Subsequently, we fix the batch size as 1 and evaluate the effect of sequence lengths on model efficiency. Specifically, we compare the model performance from two aspects: prompt length and generation length. For convenience, we sample 100 examples from the GSM8K dataset for evaluation. We conduct both 0-shot and 5-shot experiments, corresponding to short and long prompt lengths, respectively. For each experiment setting, we evaluate different generation lengths, considering values of 64, 128, 256, 512, 1024, and 2048 tokens. Since AR cannot precisely control their effective generation length, we enforce them to generate an <eos> token until the generation exceeds the target length for fair comparison. We adopt a similar approach for block diffusion.
@@ 5.2 RQ2: How can we theoretically explain the variations in inference throughput?
For this question, we propose a unified theoretical perspective for explaining variations in inference throughput. We first decompose throughput into hardware-side computational performance (FLOPs/s, determined by the roofline model and arithmetic intensity) and model-side computational efficiency (FLOPs/token). Based on this formulation, we establish asymptotic expressions of FLOPs, memory operations (MOPs), arithmetic intensity and FLOPs per token for AR, diffusion, and block diffusion models during decoding.
Using the roofline model, we further analyze the decoding behaviors of the three architectures—autoregressive, diffusion, and block diffusion models—under varying sequence lengths and batch sizes. Together, these analyses provide a unified theoretical basis for understanding the throughput variations observed in Section  5.1.
Table 1: Asymptotic FLOPs, Memory Operations (MOPs), and Arithmetic Intensity (ArInt) for autoregressive (AR), block diffusion, and diffusion language models (DLMs). Here $L=L_{p}+L_{g}$ (prompt + generated length), $B$ is batch size, $d$ is hidden dimension, $K$ is the number of diffusion steps, and $G$ is the block size. The first and second terms in each expression correspond to feed-forward and attention operations, respectively.
@@
FLOPs $\mathcal{O}(BL_{g}d^{2})+\mathcal{O}(BL_{g}Ld)$ $\mathcal{O}(KBGd^{2})+\mathcal{O}(KBGLd)$ $\mathcal{O}(KBLd^{2})+\mathcal{O}(KBL^{2}d)$
@@
ArInt $\approx\begin{cases}\mathcal{O}(B),&L_{p}\ll d\\ \mathcal{O}(1),&L_{p}\gg d\end{cases}$ $\approx\begin{cases}\mathcal{O}(BG),&L_{p}\ll d\\ \mathcal{O}(G),&L_{p}\gg d\end{cases}$ $\approx\begin{cases}\mathcal{O}(BL),&L_{p}\ll d\\ \mathcal{O}(L),&L_{p}\gg d\end{cases}$
@@
Figure 3: Roofline comparison of varying decoding architectures (AR, DLM, and Block Diffusion). The simulation uses the NVIDIA RTX A800 80 G platform, with a peak performance of 312 TFLOP/s and an arithmetic intensity ridge point of 153 FLOPs/Byte.
5.2.1 Theoretical Analysis of Inference Throughput
To interpret the throughput variations observed in RQ1, we introduce a unified analytical perspective that decomposes inference throughput into two components: (1) hardware-side computational performance, representing the amount of computation effectively executed per unit time (FLOPs/s) and determined by the roofline model and the workload’s arithmetic intensity; and (2) model-side computational efficiency, defined as the average computational cost required to generate a single token (FLOPs/token).
Accordingly, the inference throughput can be formally expressed as:
$\text{Throughput}=\frac{\text{Generated Tokens}}{\text{Generation Time(seconds)}}=\frac{\text{FLOPs~per~second}}{\text{FLOPs ~per~token}}.$ (3)
Hardware-Side Performance. The roofline model characterizes the upper bound of attainable computational performance under given hardware constraints. And arithmetic intensity, defined as the ratio between computation and memory access, determines whether a workload is memory-bound (limited by bandwidth) or compute-bound (limited by the peak FLOP performance). When arithmetic intensity is below the ridge point, the effective compute is limited by memory bandwidth; when above, it reaches the compute bound determined by the peak FLOPs.
Model-Side Efficiency. The average FLOPs per token is a statistical measure of the model’s computational efficiency. It is obtained by dividing the model’s theoretical asymptotic FLOPs—representing the total floating-point operations required for the entire generation process—by the total number of generated tokens.
@@ 5.2.1 Theoretical Analysis of Inference Throughput
Architectural Comparison. We consider three decoding architectures—AR, diffusion, and block diffusion—and derive their asymptotic FLOPs, memory operations (MOPs), arithmetic intensity, and FLOPs per token during decoding. Table 1 summarizes these asymptotic expressions for each architecture. To provide a hardware reference, we also include the roofline model of our evaluation platform—as shown in Figure 3. Under this reference, we analyze the decoding behavior of the three architectures at a batch size of 1. AR models lie near the bandwidth roof and are therefore memory-bound. Diffusion models generally fall within the compute-bound regime, even for short sequences, because full-sequence denoising results in high arithmetic intensity and rapidly increasing FLOPs with sequence length. Block diffusion models occupy an intermediate region: still memory-bound, but with higher arithmetic intensity than AR due to parallel block-wise generation that increases computation per memory access.

### v3

@@ Abstract
Diffusion language models (DLMs) have emerged as a promising alternative to the long-dominant autoregressive (AR) paradigm, offering a parallelable decoding process that could yield greater efficiency. Yet, in practice, current open-source DLMs often underperform their AR counterparts in speed, limiting their real-world utility. This work presents a systematic study of DLM efficiency, identifying key issues in prior evaluation methods. Through empirical benchmarking and a theoretical analysis, we demonstrate that AR models generally achieve higher throughput, while DLMs consistently lag. We also investigate acceleration strategies, finding that techniques like dual cache and parallel decoding mainly offer gains at small batch sizes, with their benefits diminishing upon scaling. Our findings underscore the necessity of robust evaluation methods and improved acceleration strategies to advance research on DLMs.
@@ 1 Introduction
Diffusion language models (DLMs) have recently emerged as a competitive alternative to the long-dominant autoregressive (AR) approach [llmsurvey-arxiv-2023-wayne] for natural language generation. Despite the relatively low profile of non-autoregressive research, its persistent advancement has now yielded performant models like LLaDA [dllm-arxiv-2025-shen] that are competitive with their AR counterparts. This represents a significant departure from the sequential generation paradigm that has defined the field for years.
[…]
@@
  • Evaluation Scope: Prior evaluations often compare DLMs and AR models in limited conditions, such as single instance inference or fixed generation length. These constrained conditions fail to capture performance [^1] In this paper, performance refers to system-level efficiency metrics rather than task-level or downstream accuracy. across diverse settings and poorly represent real-world use.
@@
  • Throughput modeling and theoretical analysis: We conduct a theoretical analysis aiming to provide a fine-grained understanding of inference throughput. Specifically, we model throughput as a function of hardware-side performance (FLOPs/s) and model-side efficiency (FLOPs/token), and analytically characterize how autoregressive, diffusion, and block diffusion models behave under varying sequence lengths and batch sizes.
@@ 2.1.3 Block Diffusion Language Models
Block diffusion models combines autoregressive dependencies across blocks with parallel diffusion refinement within each block, maintaining long-range contextual coherence while enabling efficient parallel generation. This architecture also naturally supports KV caching, enhancing inference throughput. It introduces a complementary attention mask that enables bidirectional attention within blocks while preserving autoregressive dependencies across them. During inference, this design allows the model to perform autoregressive generation across blocks and parallel diffusion refinement within each block, leading to a significant reduction in training cost and improved inference efficiency. For example, BD3-LM [5-iclr-2025-marianne], trained from scratch, demonstrates that block-level diffusion can substantially improve efficiency without sacrificing generation quality. Building on this direction, D2F and Fast-dLLM v2 [6-arxiv-2025-chengyue] show that block diffusion-style generation can also be achieved by training from existing models—D2F from diffusion backbones and Fast-dLLM v2 from autoregressive ones.
@@ 2.2.1 Reducing Per-step Computational Overhead
Sequence-level KV Cache. One line of works store and reuse key/value states across decoding steps: The conventional KV cache is designed for strictly autoregressive decoding and does not directly apply to DLMs. To address this, recent work shows that redesigning the decoding schedule allows DLMs to recover much of the efficiency benefit of KV caching. For example, Fast-dLLM [7-arxiv-2025-chengyue] introduces a DualCache (prefix + suffix) to improve reuse while bounding quality loss. DPad [8-arxiv-2025-xinhua] proposes a suffix-window and distance-decay dropout to restrict attention and thereby limit suffix token caching and computation. And Sparse‑dLLM [9-arxiv-2025-yuerong] develops a training-free dynamic cache eviction framework that retains only salient token states in cache and evicts less relevant prefix/suffix entries to improve throughput.
Step-level Feature Reuse. Another line of works reuse stable intermediate representations inside each denoising iteration. dKV-Cache [10-arxiv-2025-xinyin] introduces a delayed caching scheme, where a token’s key and value states are cached one denoising step after it is decoded. dLLM-Cache [21-arxiv-2025-zhiyuan] uses feature similarity to identify stable tokens and reuse their cached features., and FreeCache [11-arxiv-2025-zhanqiu] reuses stable KV states of early clean tokens while updating only actively changing ones. These methods are complementary to sequence-level caching.
@@ 2.2.2 Reducing Diffusion Sampling Steps
Step Reduction via Distillation. A key direction for accelerating DLMs is reducing the number of denoising steps through distillation. Early works on diffusion models, such as Progressive Distillation [12-iclr-2022-tim], adopt teacher–student schemes that iteratively halve the required sampling steps. Di4C [13-arxiv-2024-satoshi] extends this approach to discrete diffusion by explicitly distilling inter-token correlations. Recently, DLM-One [14-arxiv-2025-tianqi] apply these distillation ideas to continuous DLMs, using score-based distillation to train a one-step generator that produces the full sequence in a single forward pass, achieving substantial speedups while maintaining near-teacher quality.
@@ 3 Efficiency Evaluation Issues
Efficiency Metrics. Efficiency reporting are inconsistent across studies, making it difficult to compare models fairly. Some studies report latency per token [11-arxiv-2025-zhanqiu], while others focus on throughput (tokens per second) on specific hardware setups [9-arxiv-2025-yuerong, 10-arxiv-2025-xinyin]. These metrics, while useful, often fail to capture the complete picture. For example, a model might show high throughput, but if each step involves heavy computation, the overall cost could still be high. This is especially relevant for DLMs, which often require more operations per token than AR models. To support clearer and more practical comparisons, it would be helpful to include standardized metrics that consider both computational cost and decoding performance. Examples might include FLOPs per token or throughput under fixed resource constraints. These indicators can offer a more balanced view of model efficiency across different architectures.
@@ 4.1 Overview and Research Questions
RQ2: How can we theoretically explain variations in inference efficiency? We aim to provide a fine-grained theoretical analysis that interprets throughput in terms of hardware-side performance (FLOPs/s) and model-side efficiency (FLOPs/token). This analysis helps explain the empirical results in RQ1, reveal key computational bottlenecks, and suggest potential directions for future acceleration.
@@ 4.2.2 Benchmark and Scenarios
We use the GSM8K dataset for all performance benchmarks. The prompt length is controlled by adjusting the number of examples (i.e., 0-shot and 5-shot) in the prompt. For generation length, DLMs allow direct parameter control. For AR and block diffusion models, we enforce generation to continue until the target length is met (i.e., ranging from 64 to 2048), even after an “<eos>” token is produced, to ensure a fair comparison of throughput.
@@ 4.2.3 Models and Acceleration Methods
Specifically for the LLaDA model, we evaluate two acceleration strategies introduced in the Fast-dLLM series: the dual cache strategy (to reduce per-step computation) and the confidence-aware parallel decoding strategy (to reduce the number of steps). For the block diffusion, which inherently supports KV cache utilization due to its block-wise autoregressive design, we analyze the confidence-aware parallel decoding strategy. All hyperparameters for these acceleration methods are configured according to their original papers’ settings for GSM8K, and we further verify that the accuracy remains aligned with the original models.
@@ 4.2.4 Evaluation Metrics
Arithmetic Intensity. In model inference, the overall latency is determined by both computational workload and memory access. The computational workload is typically measured in floating-point operations (FLOPs), while memory activity is quantified in bytes of read and write operations (MOPs). Furthermore, the arithmetic intensity (Luebke et al. [gpgpu-siggraph-2004-david]) is defined as the ratio of floating-point operations to memory operations, measuring the balance between computation and data movement:
[…]
Arithmetic intensity characterizes how many floating-point operations are executed for each byte of data moved. A low arithmetic intensity indicates that the workload requires frequent memory access relative to computation, meaning the attainable performance is limited by memory bandwidth—a scenario referred to as memory-bound. Conversely, a high arithmetic intensity implies that the operation performs many computations per byte transferred, and the performance bottleneck shifts to the GPU’s compute capacity, known as a compute-bound regime.
To understand the relationship between arithmetic intensity and performance, we adopt the roofline model [15-arxiv-2009-samue] for our analysis, following the approach of Kim et al. [16-arxiv-2025-minseo]. This model relates a system’s peak compute performance to its memory bandwidth, identifying whether a workload is compute-bound or memory-bound. In this framework, the arithmetic intensity ridge point ($\text{Arithmetic Intensity}_{\text{ridge}}$) defines the boundary between the two regimes (compute-bound and memory-bound) and is calculated as follows:
@@
According to the roofline model, workloads with arithmetic intensity below this ridge point are memory-bound regime, while those above it are compute-bound regime. In the memory-bound regime, the attainable performance $P$ scales linearly with arithmetic intensity, with a slope equal to the peak sustainable memory bandwidth $B_{\text{mem}}$:
$P=B_{\text{mem}}\times\text{Arithmetic Intensity}.$
In contrast, in the compute-bound regime, the performance reaches the flat ceiling defined by the peak floating-point performance $P_{\text{max}}$.
@@ 5.1.1 Effect of Sequence Length
Experimental Settings. We select LLaDA-8B-Instruct, LLaMA3.1-8B-Instruct, and Fast-dLLM v2-7B as the representative models for DLM, AR and block diffusion. Subsequently, we fix the batch size as 1 and evaluate the effect of sequence lengths on model efficiency. Specifically, we compare the model performance from two aspects: prompt length and generation length. For convenience, we sample 100 examples from the GSM8K dataset for evaluation. We conduct both 0-shot and 5-shot experiments, corresponding to short and long prompt lengths, respectively. For each experiment setting, we evaluate different generation lengths, considering values of 64, 128, 256, 512, 1024, and 2048 tokens. Since AR cannot precisely control their effective generation length, we enforce them to generate an “<eos>” token until the generation exceeds the target length for fair comparison. We adopt a similar approach for block diffusion.
@@ 5.2 RQ2: How can we theoretically explain the variations in inference throughput?
To better understand the throughput variations observed in RQ1, we conduct a theoretical analysis of how inference throughput varies with critical factors (i.e., sequence length and batch size) across different model architectures. In a prior study, Kim et al. [16-arxiv-2025-minseo] adopt the roofline model [15-arxiv-2009-samue] to derive asymptotic formulations of FLOPs, MOPs, and arithmetic intensity for AR and DLMs, thereby revealing whether each model operates in a compute- or memory-bound regime. This analysis provides valuable intuition about hardware utilization. Inspired by their work, we focus on the overall inference throughput and develop an extended analytical framework that explicitly connects model-side computational efficiency (FLOPs per token) with hardware-side computational performance (FLOPs per second). This formulation allows us to interpret throughput variations through both hardware and model-architecture perspectives, revealing how architectural design, sequence length, and batch size affect inference efficiency.
5.2.1 Theoretical Analysis of Inference Throughput
To formalize how inference throughput is influenced by hardware performance and model design, we decompose throughput into two components: (1) hardware-side computational performance, representing the amount of computation effectively executed per unit time (FLOPs/s) and determined by the workload’s arithmetic intensity and the roofline model; and (2) model-side computational efficiency, defined as the average computational cost required to generate a single token (FLOPs/token). Formally, the inference throughput can be expressed as the ratio between these two components, as shown below:
[…]
(a) ArInt (0-shot).
(b) ArInt (5-shot).
(c) ArInt (long generation).
Figure 3: (a) and (b) show the arithmetic intensity comparison across different generation lengths under the 0-shot and 5-shot settings, respectively, while (c) presents the arithmetic intensity comparison between AR and block diffusion models under longer generation lengths.
(a) ArInt (0-shot).
(b) ArInt (5-shot).
Figure 4: The arithmetic intensity comparison across different batch sizes under the 0-shot (a) and 5-shot (b) settings, respectively.
Hardware-Side Performance. To estimate the hardware-side attainable performance, we first compute the arithmetic intensity of each model and then apply the roofline model to obtain its attainable FLOPs per second. Specifically, based on the formulations proposed by Kim et al. [16-arxiv-2025-minseo], we derive simplified asymptotic expressions for FLOPs per second under both memory- and compute-bound conditions, as summarized in Table 1 (rows of FLOPs per second). Building on the formulation in Table 1, we further perform detailed approximations of arithmetic intensity under the practical experimental settings used in RQ1, with the corresponding formulas provided in Appendix A. Figure 3 and Figure 4 present these estimated values together with the hardware-specific arithmetic intensity ridge, which separates the memory-bound and compute-bound regimes. Configurations below the ridge are approximately memory-bound, whereas those above are compute-bound. This analysis allows us to infer the attainable computational performance for each setting, providing the basis for the subsequent throughput interpretation.
Table 1: Analysis of FLOPs per second and FLOPs per token for AR, block diffusion, and DLMs. The gray rows are derived based on the asymptotic analysis proposed by Kim et al. [16-arxiv-2025-minseo], following their approximate formulations and notations. Here, $B_{\text{mem}}$ denotes the peak sustainable memory bandwidth, $P_{\text{max}}$ is the peak floating-point performance, $L=L_{p}+L_{g}$ (the total length of prompt and generated response), $B$ is batch size, $d$ is hidden dimension, $K$ is the number of diffusion steps, and $G$ is the block size.
@@ 5.2.1 Theoretical Analysis of Inference Throughput
FLOPs per second (compute-bound) $P_{\text{max}}$ $P_{\text{max}}$ $P_{\text{max}}$
Generated Tokens $BL_{g}$ $BL_{g}\approx BK$ $BL_{g}\approx BK$
@@
Model-Side Efficiency. As shown in Equation 4, the average FLOPs per token is a statistical measure of the model’s computational efficiency. It is obtained by dividing the model’s theoretical asymptotic FLOPs—representing the total floating-point operations required for the entire generation process—by the total number of generated tokens. The results of the three architectural model types are shown in Table 1.